A building casts a 20ft shadow at the same time a person who is 5'3 casts a 10ft shadow. How tall is the building?
Answer by mananth(16946)

This is based on similarity of triangles
x/20 = 5' 3" /10
x= (20*21/4 )/10
x= 5*21/10
x=10.5 feet
10' 6"